OpenStudy (anonymous):

Elissa can do a job in 5 hours. Elissa and Lori working together can do the same job in 2 hours. How long would it take Lori to do the job alone?

OpenStudy (anonymous):

i got 4 2/3 hours

OpenStudy (anonymous):

Rate of Elissa is 1/5 ; let the rate of Lori be 1/x . Now 1(1/5) + 1/(1/x) = 2 ; find x .

OpenStudy (anonymous):

It should be 1/(1/5)

OpenStudy (anonymous):

that doesnt make sense. I tried again and got 2 2/3 hours

OpenStudy (anonymous):

helpp?!!

OpenStudy (dumbcow):

you are solving \[\frac{1}{5} + x = \frac{1}{2}\] \[x = \frac{1}{2} - \frac{1}{5} = \frac{5}{10} - \frac{2}{10} = \frac{3}{10}\] the rate is 3/10 therefore time is 10/3

OpenStudy (anonymous):

3 1/3 hours?

OpenStudy (whpalmer4):

Elissa's rate (\(R_E\)) is 1 job / 5 hours, or 1/5 jobs/hour. Elissa + Lori working together have a rate (\(R_E+R_L\)) of 1 job /2 hours, or 1/2 jobs/hour. You want to find Lori's rate (\(R_L\)). Standard rate formula is \(w = rt\) where \(w\) is work done, \(r\) is rate, \(t\) is time. For 1 job, \(w=1\). \(r = R_E+R_L\). \(t = 2\). \[1 = (R_E+R_L)(2)\]Substitute known value for \(R_E\):\[1=(\frac{1}{5}+R_L)(2)\]Multiply through by 5 \[5=2+10R_L\]\[3=10R_L\]\[R_L = 3/10\] So, for Lori to do it alone: \[1 = R_L*t\]\[1=(3/10)t\]\[t=10/3 = 3\frac{1}{3}\]

OpenStudy (anonymous):

@oholmes : Focus on the word "alone" , not together .

OpenStudy (whpalmer4):

Checking the work: if Lori takes 3 1/3 hours to do the whole job herself, in 2 hours she will accomplish \[\frac{2}{3 \frac{1}{3}} = 6/10\] of the job. If Elissa takes 1 hour to do 1/5 of a job, in 2 hours she will do 2/5 of the job. In two hours together, they will have done 6/10 + 2/5 = 6/10 + 4/10 = 10/10 or all of the job. The answer is correct.

OpenStudy (whpalmer4):

Figure out the individual rate for each worker, pump, drain, machine, etc. Then combine them as specified by the problem statement and you should get the right answer.